the sum of of 4 times a number, and -2, is equal to the sum of 5 times a number, and -2. Find the number
Answer by jndarrell(58)

The word "Sum" = addition.
"A number" = the unknown, aka 'x'.
"Is equal to" is =.
Go piece by piece. I will set your equation exactly as it is worded as you read it. Follow along:
4x + (-2) = 5x + (-2)
Imagine there is a "1" infront of the parenthesis around the -2. A positive * a negative always = negative. So....
4x-2=5x-2
Get x alone on one side of the equation. I am going to chose to subtract 4x from the 5x. Resulting in:
-2 = x-2
X is still not alone. Now I will add the -2 to the other side, accomplishing x alone.
0=x or x=0
To check the answer, take the found value for x and plug it back into the original equation. Anything multiplied by 0 will equal 0. You are left with -2 =-2, which is true.
:)